OpenStudy (anonymous):

Consider the compound Sn(CO3)2 and answer the following questions. What is the name of this compound? What is the molar mass of this compound? Explain how you calculated this value THIS IS MY LAST QUESTION ON MY 45 QUESTION EXAM CAN SOME ONE PLEASE HELP..

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

tin carbonate Molar mass is found by adding atomic masses for all the atoms in the molecule and expressing the amount with units of grams. Sn 119g C 12g * 2 = 24 g O 16g * 6 = 96 g 119g + 24g + 96g = 239g/mol